Ticket Number: HEA-NOC/20000331-02 Ticket Status: OPEN
Ticket Type: Scheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: listserv
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: 20000331 15:53:21 UTC Problem Start: 20000331 16:44
Ticket Closed: Problem End:
Digital/Mentec engineer to install new backup DLT tape drive on Tuesday April 4 at 15:00 for two hours.
All mail to listserv will be affected and will backup at host mail smtp server
until listserv restored.
Time to Fix:
Approx 2 hours